OpenStudy (anonymous):

a pair of sneakers cost75$. the sales tax is 4%what is the total cost

OpenStudy (anonymous):

75+ 4%of 75 =75+ 4*75/100 =75+ 3=78

OpenStudy (anonymous):

75x.04=3

OpenStudy (anonymous):

no , 75* 0.04 =3

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

So 75+3=78

OpenStudy (anonymous):

can not be higher thain thain the original cost

OpenStudy (anonymous):

you are adding sales cost

OpenStudy (anonymous):

oooooo